OpenStudy (devonhoward15):

which point is on the graph of y=4/X A.1/4 B.2/8 C.2,1/2 D.1/4,2

OpenStudy (fortytherapper):

A point has 2 values. 2 of these answers only have 1 value

OpenStudy (devonhoward15):

So either a or b?

OpenStudy (devonhoward15):

cant be c or d because they have 2 anwsers?

OpenStudy (fortytherapper):

What you want is two values, since its asking for a value (X, Y)

OpenStudy (mathmale):

It would likely profit you, as wello as the rest of us, if you'd show your actual work. For example, in (d), x= 1/4 and y = 2. Substitute these values into y=4/x and determine whether the equation is then true or false. If true, (1/4, 2) is a solution; if not (1/4, 2) is not a solution.;

OpenStudy (mathmale):

Would you please be good enough to inform those helping you if you plan to go off line.

OpenStudy (fortytherapper):

^ Agreed. People that just want direct answers make me want to eat healthy food